# Cosmology on the Generalized Proca Theory

### v1

@@ 1.4.1 Solution to the Flatness Problem
Figure 1.6: Evolution of the logarithm of the total density parameter $\log{\Omega_{\text{total}}}$. The first part of the curve is an arbitrary initial condition, the inflation process eliminates these conditions and makes the Universe highly flat. $\Omega_{\text{total}}\sim 1$, yet, in the future, the Universe could tend to deviate from 1 as observed at the end of the curve, taken from [52].
1.4.2 Solution to the Horizon Problem
At the beginning of the Universe, all the energy content was closer together. Due to accelerated expansion, the causal zone of our observable Universe is smaller today. Accelerated expansion allows thermal equilibrium between causally disconnected extremes today because before inflation they were connected and could reach thermal equilibrium. Once inflation ends, the causal connection between the extremes is lost, reducing the causal zone of our current Universe compared to the young Universe, as schematized in Figure 1.7. Additionally, to achieve the initial conditions today and properly maintain all processes during the evolution of the Universe (Nucleosynthesis, CMB, structure formation, accelerated expansion due to dark energy), the inflation process must be perfectly exponential and occur in the first $10^{-34}s$ or during 60 $e$-folds so that the Universe expands by a factor of $a_{\text{final}}/a_{\text{inicial}}\sim 10^{43}$[52, 59].
Figure 1.7: Schematic of the change in the size of the Universe due to the inflation process, taken from [52].
@@ 1.5 Observational Constraints and Dark Energy
As has been discussed, the Universe today is experiencing accelerated expansion [48, 4]. Similar to inflation, this accelerated expansion can be explained by the fluid known as "dark energy," whose main characteristic is that its equation of state satisfies $w<-1/3$, meaning it can produce accelerated expansion. The first attempt to introduce dark energy in cosmological models was $\Lambda$. However, due to previously mentioned issues, such as the nature of $\Lambda$, the Hubble tension [5, 6, 1, 7], the $S_{8}$ tension [8, 9, 10, 11], anomalies in the cosmic microwave background (CMB) [12, 13], and the cosmic dipole misalignment [14, 15, 16], among others [17, 18, 19]. It is convenient to describe dark energy with additional degrees of freedom. These can include tachyonic fields [20, 60], quintessence fields [20], vector fields [61, 62, 63, 64, 65, 66, 67], p-forms [68, 24, 69], among other possible descriptions that can generate a viable model of dark energy.
@@
Figure 1.8: Representation of the history of the Universe as a function of time and temperature, indicating the type of dominance in each epoch. Furthermore, the first seconds of the Universe are still not well understood, taken from [47].
@@ 1.6 Dynamical System Approach
One of the most effective tools for analyzing the parameter space of theoretical models is the dynamical systems approach [20, 41, 42, 70]. This technique is particularly powerful in cosmology, where the equations of motion for the fields are reformulated using dimensionless variables, typically derived from the first Friedman equation. This transformation results in a set of autonomous first-order differential equations, with their stationary solutions residing in the phase space of the dynamical variables.
@@
where $A^{i}$ is a constant depending on the initial conditions. Depending on the sign of $\text{Re}(\lambda_{i})$, the fixed point can be classified as an attractor or a repeller in the direction of $e_{u^{i}}$, such that:
@@ 1.6.1 Numerical approach
These constraints ensure that the numerical solutions of the autonomous equations align with physically meaningful properties of the model. For instance, imposing the condition $w_{\text{eff}}<-1/3$, where $w_{\text{eff}}$ represents the effective equation of state, ensures that any resulting solution describes an accelerated expansion of the universe. Other constraints may include non-negativity of density parameters for consistency with physical laws. Solutions meeting these criteria are labeled as “cosmologically viable,” while those failing to satisfy them are categorized as “non-viable” and excluded. It is important to note that all cosmologically viable solutions derived from the same parameter set are indistinguishable in terms of physical interpretation.
@@
A large set of $N$ random points in the parameter space $\{a_{1},\ldots,a_{k}\}$ is generated. For each set of parameters, fixed points are determined by numerically solving the system of algebraic equations $\{f_{i}(x_{1},\ldots,x_{n})=0\ |\ i=1,\ldots,n\}$ [^4] This can be achieved using standard numerical methods, such as NSolve in Mathematica.. Subsequently, the physical constraints from the prior step are applied to classify the solutions. Points that fulfill the constraints are marked as “cosmologically viable,” whereas those that do not are labeled “non-viable” and removed from further consideration. For models involving two parameters, this stochastic exploration may result in outcomes similar to those illustrated schematically in Figure 1.9. In this context, all cosmologically viable solutions correspond to accelerated states of the system, such as Dark Energy (DE)-dominated fixed points.
@@
Here, we are mainly interested in DE dominated attractor points. Therefore, once an attractor is found, the computation of the Jacobian matrices for the given set of parameters is stopped, just to proceed to another point in the parameter space.
In the chapter 2, we will illustrate the numerical method described here by analyzing the asymptotic behavior of a specific model, prioritizing the search of DE domination points.
1.7 Example: Dark energy model based on a tachyon field.
Tachyonic fields essentially originate from the D-brane theory in string theory, where the tachyonic field arises from the excitation of spacetime instabilities (commonly used to describe density perturbations or the reheating stage after inflation). These fields are characterized by having a complex mass or being superluminal [^5] A field or particle is said to be superluminal when its propagation speed exceeds the speed of light.[20], such that the contribution of the tachyonic field is described by a D3-brane as
$S_{\phi}=-\int\text{d}^{4}xV(\phi)\sqrt{-\det{(g_{\mu\nu}+\partial_{\mu}\phi\partial_{\nu}\phi)}},$ (1.43)
where $V(\phi)$ represents the tachyonic potential and $\phi$ denotes the tachyonic field. The functional form of the potential $V$ can vary depending on the phenomena being described. However, the specific details of the potential are beyond the scope of this thesis. Instead, the focus lies on the cosmological effects it may produce. Notably, for the model to be a viable candidate for dark energy, it is crucial that $\partial_{\mu}\phi\ll 1$, as will be demonstrated below:
When the condition $\partial_{\mu}\phi\ll 1$ is satisfied, the determinant can be approximated as
$\det{(g_{\mu\nu}+\partial_{\mu}\phi\partial_{\nu}\phi)}\approx g+gg^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi,$ (1.44)
which implies that the effective action is expressed in terms of a field that, while not superluminal, originates from the same functional structure, given by
$S_{\phi}=-\int\text{d}^{4}xV(\phi)\sqrt{-g}\sqrt{1+\partial^{\mu}\phi\partial_{\mu}\phi}.$ (1.45)
Thus, by varying the action with respect to the metric, the energy-momentum tensor associated with the tachyonic field is obtained as, [^6] The general energy momentum tensor can be computed using the expression $T_{\mu\nu}=g_{\mu\nu}\mathcal{L}-2\frac{\delta\mathcal{L}}{\delta g^{\mu\nu}}$.
$T_{\mu\nu}=\frac{V(\phi)\partial_{\mu}\phi\partial_{\nu}\phi}{\sqrt{1+g^{\alpha\beta}\partial_{\alpha}\phi\partial_{\beta}\phi}}-g_{\mu\nu}V(\phi)\sqrt{1+g^{\alpha\beta}\partial_{\alpha}\phi\partial_{\beta}\phi}.$ (1.46)
Assuming that the geometry is described by a flat FLRW metric ($k=0$) as shown in Eq. (1.12), we choose the field to depend solely on time, i.e., $\phi\equiv\phi(t)$, in order to maintain the same symmetry as the metric for simplicity. As a result, the energy density $\rho_{\phi}$ and pressure $p_{\phi}$ associated with the field are expressed as:
$\rho_{\phi}=-T_{0}^{0}=\frac{V(\phi)}{\sqrt{1-\dot{\phi}^{2}}},$ (1.47)
$p_{\phi}=T_{i}^{i}=-V(\phi)\sqrt{1-\dot{\phi}^{2}},$ (1.48)
respectively. Therefore, the equation of state for this dark energy model becomes
$w_{\text{DE}}=-1+\dot{\phi}^{2},$ (1.49)
implying that for the tachyonic field to drive accelerated expansion, it is necessary that $|\dot{\phi}|<\sqrt{2/3}$. In the limit $\dot{\phi}\sim 0$, the cosmological constant effect is recovered. However, in general, the equation of state for the tachyonic field can evolve dynamically throughout the history of the universe, as $\dot{\phi}$ is not necessarily constant. This makes it a model for dynamical dark energy.
1.7.1 Dynamical System
The action encompassing all cosmological effects, including geometry, the tachyonic fluid, and perfect fluids of radiation and matter, is given by
$S\equiv\int\text{d}^{4}x\sqrt{-g}\left(\frac{m_{\text{P}}^{2}R}{2}-V(\phi)\sqrt{1+\partial_{\mu}\phi\partial^{\mu}\phi}+\mathcal{L}_{m}+\mathcal{L}_{r}\right),$ (1.50)
Then, the Friedman equations are expressed as
$1=\frac{\rho_{\phi}}{3m_{\text{P}}^{2}H^{2}}+\frac{\rho_{r}}{3m_{\text{P}}^{2}H^{2}}+\frac{\rho_{m}}{3m_{\text{P}}^{2}H^{2}},$ (1.51)
$-2m_{\text{P}}^{2}\dot{H}=\dot{\phi}^{2}\rho_{\phi}+\frac{4}{3}\rho_{r}+\rho_{m},$ (1.52)
and the equation of motion for the tachyon field is
$\frac{\ddot{\phi}}{1-\dot{\phi}^{2}}+3H\dot{\phi}+\frac{1}{V}\frac{\mathrm{d}V}{\mathrm{d}\phi}=0.$ (1.53)
The continuity equations for radiation and matter are given by Eqs. (1.24) and (1.26), respectively [60, 20].
Before solving the equations describing the dynamics of the universe, it is necessary to introduce dimensionless variables to establish an autonomous system of differential equations. For this model, the standard variables are defined as [60]:
$x=\dot{\phi},\quad y^{2}=\frac{V}{3m_{\text{P}}^{2}H^{2}},\quad\Omega_{m}=\frac{\rho_{m}}{3m_{\text{P}}^{2}H^{2}},\quad\Omega_{r}=\frac{\rho_{r}}{3m_{\text{P}}^{2}H^{2}}.$ (1.54)
The Friedman equations can then be rewritten to provide a constraint and the deceleration parameter in terms of these dynamical variables:
$1=\frac{y^{2}}{\sqrt{1-x^{2}}}+\Omega_{r}+\Omega_{m},$ (1.55)
$q=\frac{1}{2}\left(1-3\sqrt{1-x^{2}}y^{2}+\Omega_{r}\right),$ (1.56)
where Eq. (1.55) allows $\Omega_{m}$ to be expressed in terms of the other variables.
1.7.2 Autonomous System
In order to obtain the equations, each of the variables can be differentiated with respect to cosmic time and the equations of motion can be used to simplify the expressions. However, if each of the dynamical variables (Eq. 1.54) is differentiated with respect to cosmic time, the resulting system is closed but all terms are proportional to the Hubble parameter $H$. In other words, the form of the equations becomes $\dot{x}=Hf(x,y,\Omega_{m},\Omega_{r})$. Therefore, it is advantageous to switch from using cosmic time to the number of e-folds $N$, where the differential relation is given by $dN=Hdt$. As a result, the system of differential equations takes the following form [^7] Here, a prime denotes differentiation with respect to the number of e-folds $N$.
$x^{\prime}$ $=$ $-\sqrt{3}\left(1-x^{2}\right)\left[\sqrt{3}x+\alpha y\right],$ (1.57)
$y^{\prime}$ $=$ $y\left[\frac{\sqrt{3}}{2}\alpha yx+q+1\right],$ (1.58)
$\Omega_{r}^{\prime}$ $=$ $2\Omega_{r}(q-1),$ (1.59)
where the potential $V$ is chosen such that it satisfies the following differential equation:
$\alpha=m_{p}\frac{V_{,\phi}}{V^{3/2}},$ (1.60)
with $\alpha$ being a constant.
1.7.3 Dynamical Analysis
Before solving the differential equations, it is essential to study their behavior since they are parametrically dependent on $\alpha$. This allows for finding the fixed points and analyzing their stability. From these, the values of $\alpha$ can be conditioned to match observational constraints. Additionally, the universe must pass through an era dominated by radiation, characterized by having the effective equation of state $w_{\text{eff}}\equiv(2q-1)/3=1/3$. Its stability corresponds to a saddle point because the universe must pass through it but does not remain at this point. It should then transition to a matter-dominated era characterized by $w_{\text{eff}}=0$, where the stability is also a saddle point, similar to radiation. Finally, the universe must enter a dark energy-dominated era characterized by $w_{\text{eff}}<-1/3$, where stability is attractive, as dark energy tends to dominate the evolution once this era is approached.
Fixed Points
  • ($R$) Radiation Dominance
    $x=0,\quad y=0,\quad\Omega_{r}=1,\quad$ (1.61)
    with $\Omega_{\text{DE}}\equiv\frac{\rho_{\phi}}{3m_{\text{P}}^{2}H^{2}}=0$, $w_{\text{DE}}=-1$, $\Omega_{m}=0$, and $w_{\text{eff}}=1/3$.
  • ($M$) Matter Dominance
    $x=0,\quad y=0,\quad\Omega_{r}=0,\quad$ (1.62)
    with $\Omega_{\text{DE}}=0$, $w_{\text{DE}}=-1$, $\Omega_{m}=1$, and $w_{\text{eff}}=0$.
  • ($DE$) Dark Energy Dominance
    $x=\mp\frac{\alpha\sqrt{-\alpha^{2}+\sqrt{36+\alpha^{4}}}}{3\sqrt{2}},\quad y=\pm\frac{\sqrt{-\alpha^{2}+\sqrt{36+\alpha^{4}}}}{\sqrt{6}},\quad\Omega_{r}=0$
    with $\Omega_{\text{DE}}=1$, $\Omega_{m}=0$, and
    $w_{\text{DE}}=w_{\text{eff}}=-1+\frac{1}{18}\alpha^{2}\left(-\alpha^{2}+\sqrt{36+\alpha^{4}}\right),$ (1.63)
    where the interest lies in ensuring that dark energy generates accelerated expansion, i.e., $w_{\text{DE}}<-1/3$. This condition is satisfied when $|\alpha|<\sqrt{2}3^{1/4}$.
Stability of the Fixed Points
To study the stability, it is necessary to analyze the signs of the eigenvalues of the Jacobian matrix. As we mentioned previously, If all the eigenvalues are negative, the fixed point tends to be an attractor; if all are positive, it tends to be a repeller; and if there are both positive and negative eigenvalues, it is a saddle point, attracting along the direction where the eigenvalue is negative.
  • ($R$) Radiation Dominance
    $\lambda_{1}=-3,\quad\lambda_{2}=2,\quad\lambda_{3}=1$ (1.64)
    This point corresponds to a saddle point because one eigenvalue is negative, and two are positive. Additionally, the eigenvector associated with $\lambda_{1}=1$ is $(0,0,1)$ in the phase space $\{x,y,\Omega_{r}\}$, indicating that this point is unstable in the $\Omega_{r}$-direction. Therefore, $\Omega_{r}$ will tend to move away from the value at this fixed point.
  • ($M$) Matter Dominance
    $\lambda_{1}=-3,\quad\lambda_{2}=\frac{3}{2},\quad\lambda_{3}=-1$ (1.65)
    This point corresponds to a saddle point because two of the eigenvalues are negative and one is positive. Furthermore, the eigenvector associated with $\lambda_{3}=-1$ is $(0,0,1)$, indicating that the point is stable in the $\Omega_{r}$-direction in the neighborhood of $\Omega_{r}\simeq 0$. Therefore, $\Omega_{r}\simeq 0$ is an attractor at this fixed point.
  • ($DE$) Dark Energy Dominance
    In this case, the eigenvalues have very lengthy expressions, so the analysis is done by plotting them as a function of $\alpha$.
    Figure 1.10: Eigenvalues of the Jacobian matrix as a function of $\alpha$ for the dark energy fixed point, showing that $\lambda_{1,2,3}<0$ in the region generating accelerated expansion, where $\alpha<\sqrt{2}3^{1/4}$.
    In Figure 1.10, it can be seen that, regardless of the values of $\alpha$ in the region of accelerated expansion, all eigenvalues $\lambda_{1,2,3}<0$. Therefore, the dark energy fixed point is an attractor in the parameter space.
1.7.4 Evolution example
Once the parameter space is fully classified, there are enough clues to solve the differential equations from Eq. (1.57) to (1.59), taking into account the initial conditions and observational constraints[1]. The initial conditions can be set in the radiation era with a redshift of $z_{r}\sim 2.42\times 10^{7}$. Additionally, $\alpha$ is chosen in the region where dark energy generates accelerated expansion, such that the initial conditions were:
$\begin{gathered}\alpha=0.4,\quad x_{i}=10^{-25},\quad\Omega_{r_{i}}=0.99995.\end{gathered}$ (1.66)
Since the constraint from Eq. (1.55) must be satisfied and $10<\Omega_{m}<$ 1, it follows that $|y_{i}|<0.00707$. The value that appropriately adjusted the observational conditions[1, 12, 47] was:
$y_{i}=9.02\times 10^{-14},$ (1.67)
such that $\Omega_{m_{i}}=5\times 10^{-5}$, obtaining the evolution shown in Figure 1.11.
Figure 1.11: Evolution of the density parameters, the effective equation of state of the universe, and the equation of state of dark energy throughout the expansion history of the Universe. The initial conditions, Eq. (1.66), were chosen in the early radiation era with a redshift of $z_{r}=2.42\times 10^{7}$. The universe passes through the radiation-dominated phase in the early moments (red dashed line), followed by a matter-dominated era (yellow dashed line), and ends in the dark energy-dominated era (solid black line), characterized by $w_{\text{eff}}<-1/3$ (blue dotted line). It is also observed that the equation of state of the universe matches the equation of state of dark energy (dark red dashed line) during this period.
It is important to highlight that although these tests help identify viable dark energy models within the framework of reproducing the expansion history of the Universe, they are not sufficient. This is because not all constraints have been tested. Further evaluation is required regarding possible anisotropies, the effect of this specific dynamic dark energy model on the value of $H_{0}$, and, if confirmed, a more thorough analysis in other periods of the universe to examine how it impacts structure formation, the value of $H_{0}$, and other phenomena not covered in this project.
1.8 The Generalized SU(2) Proca Theory

### v2

@@ 1.3.3 Inflationary Solution to Primordial Puzzles
Figure 1.5: Evolution of the logarithm of the total density parameter $\text{Log}\left({\Omega_{\text{total}}}\right)$. The first part of the curve is an arbitrary initial condition, the inflation process eliminates these conditions and makes the Universe highly flat. $\Omega_{\text{total}}\sim 1$, yet, in the future, the Universe could tend to deviate from 1 as observed at the end of the curve [6].
Figure 1.6: Schematic illustrating the solution to the horizon problem. A small, causally connected region before inflation (left) is expanded exponentially (center) to become vastly larger than the post-inflationary Hubble radius (right). Regions A and B, though causally disconnected today, originated from within the same causally connected pre-inflationary patch [6].
1.3.4 Remaining Challenges and Tensions
While inflation resolves key issues of the standard hot Big Bang model, the combined $\Lambda$CDM $+$ Inflation paradigm still faces significant hurdles:
The Cosmological Constant Problem
The observed value of the cosmological constant $\Lambda$, which drives the current accelerated expansion, is extremely small compared to theoretical expectations from quantum field theory. Interpreting $\Lambda$ as the vacuum energy density leads to predictions that are famously $60-120$ orders of magnitude larger than observed [45, 46].
Observational Tensions: $H_{0}$ and $S_{8}$
Recent, precise measurements have revealed statistically significant tensions between cosmological parameters inferred from early-universe probes (like the CMB by Planck) assuming $\Lambda$CDM, and those measured via late-universe probes:
  • $H_{0}$ Tension: Measurements of the Hubble constant today ($H_{0}$) using local distance ladder methods (e.g., Cepheids and SNe Ia via SH0ES) yield a value around $H_{0}\approx(73.04\pm 1.04)\,km/s\text{Mpc}$ [13, 47, 48], which is in $\sim 5\sigma$ tension with the value inferred from Planck CMB data assuming $\Lambda$CDM, $H_{0}=(67.27\pm 0.60)\,km/s\text{Mpc}$ [7, 18, 48]. This discrepancy suggests a potential breakdown in the standard model connecting the early and late universe [49, 21].
  • $S_{8}$ Tension: Measurements of the amplitude of matter clustering at late times, often parameterized by $S_{8}\equiv\sigma_{8}\sqrt{\Omega_{m}/0.3}$ (where $\sigma_{8}$ is the root mean square fluctuation amplitude on 8 Mpc scales), derived from weak lensing and galaxy clustering surveys tend to be lower than the value predicted by Planck CMB data extrapolated forward using $\Lambda$CDM. This tension is typically at the $2-3\sigma$ level and might indicate suppressed structure growth [50, 51, 48].
Figure 1.7: Schematic evolution of a comoving perfect sphere (red) in an axially symmetric Bianchi I spacetime described by Eq. (1.35). Differential expansion rates distort the sphere into an ellipsoid (blue, black) over time.
Challenges to Isotropy: Observational Anomalies
While the $\Lambda$CDM model assumes perfect large-scale isotropy and homogeneity (FLRW metric, Eq. (1.15)) and Gaussian random initial fluctuations, certain features observed primarily in the CMB maps present potential challenges or “anomalies” relative to these assumptions. These include unexpected alignments of the lowest multipole moments (quadrupole and octopole), a hemispherical power asymmetry, anomalous cold spots, and perhaps unusual patterns in polarization data [18, 19, 52]. Some studies also suggest potential large-scale anisotropies in galaxy distributions or flows [53, 54, 55].
The statistical significance and origin of these anomalies are subjects of ongoing debate – they might be residual systematic effects, foreground contamination, or simply rare statistical fluctuations within the standard model. However, their persistence motivates considering deviations from perfect isotropy. Anisotropic cosmological models, such as the Bianchi spacetimes, offer a theoretical framework for such investigations. Their general metric is described by
$ds^{2}=-dt^{2}+a^{2}(t)\gamma_{ij}dx^{i}dx^{j}.$ (1.34)
In this equation, $\gamma_{ij}=e^{2\sigma_{i}}\delta_{ij}$, and the condition $\sum_{i=1}^{3}\sigma_{i}=0$ guarantees that $a$ describes the average expansion. The Bianchi I model, for instance, describes a spatially flat but anisotropically expanding universe, whose line element can be written (in axially symmetric form) as:
$ds^{2}=-dt^{2}+a^{2}(t)\left[e^{-4\sigma(t)}dx^{2}+e^{2\sigma(t)}\left(dy^{2}+dz^{2}\right)\right],$ (1.35)
where $\sigma(t)$ quantifies the shear or deviation from isotropy (see Figure 1.7). Although CMB observations strongly constrain any residual anisotropy today ($\sigma(t_{0})\ll 1$[8], studying these models helps test the robustness of the Cosmological Principle and explore potential explanations for observed large-scale anomalies.
These outstanding problems – the theoretical puzzle of $\Lambda$, the observational tensions in $H_{0}$ and $S_{8}$, and the potential CMB anomalies – collectively provide strong motivation for exploring cosmological models beyond the standard $\Lambda$CDM paradigm.
@@ 1.4 Observational Constraints and Dark Energy
Figure 1.8: Timeline of the Universe according to the $\Lambda$CDM model including inflation. Inflation occurs at the very beginning, followed by the hot Big Bang evolution (radiation domination, matter domination, recombination/CMB emission $\sim$380,000 yrs), and late-time acceleration driven by dark energy ($\Lambda$). (Image credit: NASA/WMAP Science Team [56]).
As has been discussed, the Universe today is experiencing accelerated expansion [1, 2]. Similar to inflation, this accelerated expansion can be explained by the fluid known as dark energy, whose main characteristic is that its equation of state satisfies $w<-1/3$, meaning it can produce accelerated expansion. The first attempt to introduce dark energy in cosmological models was $\Lambda$. However, due to previously mentioned issues, such as the nature of $\Lambda$, the Hubble tension [11, 12, 7, 13], the $S_{8}$ tension [14, 15, 16, 17], anomalies in the CMB [18, 19], and the cosmic dipole misalignment [53, 54, 55], among others [20, 21, 22]. It is convenient to describe dark energy with additional degrees of freedom. These can include tachyonic fields [23, 57], quintessence fields [23], vector fields [58, 59, 60, 61, 62, 63, 64], p-forms [65, 27, 66], among other possible descriptions that can generate a viable model of dark energy.
@@ 1.5 Dynamical System Approach
Previously, we introduced the idea of defining new variables to apply the dynamical systems technique, which is useful for characterizing cosmological models. This approach is effective for analyzing the parameter space of theoretical models [23, 67, 68, 69]. This technique is particularly powerful in cosmology, where the equations of motion for the fields are reformulated using dimensionless variables, typically derived from the first Friedman equation. This transformation results in a set of autonomous first-order differential equations, with their stationary solutions residing in the phase space of the dynamical variables.
@@
where $A^{i}$ is a constant depending on the initial conditions. Depending on the sign of $\text{Re}(\lambda_{i})$, the fixed point can be classified as an attractor or a repeller in the direction of $u^{i}$ axe, such that:
@@ 1.5.1 Numerical approach
These constraints ensure that the numerical solutions of the autonomous equations align with physically meaningful properties of the model. For instance, imposing the condition $w_{\text{eff}}<-1/3$, where $w_{\text{eff}}$ represents the effective equation of state, ensures that any resulting solution describes an accelerated expansion of the universe. Other constraints may include non-negativity of density parameters for consistency with physical laws. Solutions meeting these criteria are labeled as “cosmologically viable”, while those failing to satisfy them are categorized as “non-viable” and excluded. It is important to note that all cosmologically viable solutions derived from the same parameter set are indistinguishable in terms of physical interpretation.
@@
A large set of $N$ random points in the parameter space $\{a_{1},\ldots,a_{k}\}$ is generated. For each set of parameters, fixed points are determined by numerically solving the system of algebraic equations $\{f_{i}(x_{1},\ldots,x_{n})=0\ |\ i=1,\ldots,n\}$ [^4] This can be achieved using standard numerical methods, such as NSolve in Mathematica.. Subsequently, the physical constraints from the prior step are applied to classify the solutions. Points that fulfill the constraints are marked as “cosmologically viable”, whereas those that do not are labeled “non-viable” and removed from further consideration. For models involving two parameters, this stochastic exploration may result in outcomes similar to those illustrated schematically in Figure 1.9. In this context, all cosmologically viable solutions correspond to accelerated states of the system, such as Dark Energy (DE)-dominated fixed points.
@@
Here, we are mainly interested in DE dominated attractor points. Therefore, once an attractor is found, the computation of the Jacobian matrices for the given set of parameters is stopped, just to proceed to another point in the parameter space. In the chapter 2, we will apply the numerical method to analyze the asymptotic behavior of DE model in an anisotropic background, prioritizing the search of DE domination points.
1.6 Alternative Framework: The Generalized SU(2) Proca Theory
The previously discussed challenges facing the standard $\Lambda$CDM model, including the theoretical puzzles surrounding the cosmological constant and the observational tensions in $H_{0}$ and $S_{8}$, motivate the exploration of alternative cosmological paradigms. These alternatives often involve modifying gravity itself or introducing new dynamical fields beyond the standard model components. Vector fields, particularly those associated with gauge symmetries like SU(2), represent an intriguing possibility, potentially arising from more fundamental physics frameworks and offering richer dynamics than simple scalar fields.
One such well-motivated framework is the Generalized SU(2) Proca (GSU2P) theory. This theory describes the dynamics of interacting, massive, non-Abelian vector fields ($B^{a}_{\mu}$) belonging to the Lie algebra of the SU(2) group, coupled to gravity. As detailed in Refs. [41, 78] (building on earlier work [79, 80] and extended in [81]), the action for GSU2P includes specific derivative self-interactions and couplings to curvature designed to maintain the global SU(2) symmetry and, crucially, to ensure the theory propagates the correct number of degrees of freedom (avoiding ghostly Ostrogradski instabilities often associated with higher-derivative theories [82, 83, 84, 85], although see [86, 87] for further discussion). These vector fields can potentially act as a source of dark energy or modify gravity, offering new avenues to address cosmological puzzles [38, 88]. The following section details the specific structure of the GSU2P action.